OpenStudy (ellie202000):

Find the solutions to the system. Type Equation Remarks Linear 3x + 5y = -10 A line with intercepts (-, 0) and (0, -2). Exponential y = -6(3)x This function lies in quadrants III and IV.

OpenStudy (ellie202000):

Here is the graph they came up with. How did they get the graph and the coordinates? The answers are (-3, -0.2) and (-1.6, -1). How did they come up with that?

OpenStudy (solomonzelman):

the second function should be, \(\large\color{black}{ \displaystyle y=-6(3)^x }\). the graph they posted is correct. remark - should include the second quadrant as well.

OpenStudy (solomonzelman):

AT FIRST do you understand how to graph the line, or not?

OpenStudy (ellie202000):

No that's what confusing me the most.

OpenStudy (solomonzelman):

put in a y-intercept form y=mx+b and plug in any number for x to find the y. (write your answer in terms of a coordinate (x,y) and do this again to obtain at least 2 (x,y) coordinates.) Connect the coordinates and graph your line through those coordinates/points using a ruler.

OpenStudy (ellie202000):

They say I am allowed to use a graphing calculator but I don't know how to use a graphing calculator.

OpenStudy (ellie202000):

But what if you have a problem with an exponent? Like this one:\[Y=-5^{-X}\]\[Y=X+5\]

OpenStudy (solomonzelman):

https://www.desmos.com/calculator